OpenStudy (anonymous):

I need help writing a word problem. Write a word problem describing the time it takes to complete an activity individually and with a friend. For example, if John takes 2 hours to mow his lawn and it takes his sister Maria 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together?

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

what is wrong with that example?

OpenStudy (anonymous):

I can't use it

OpenStudy (anonymous):

you need to come up with a different one?

OpenStudy (anonymous):

smith can paint a room in 4 hours wesson takes 5 hours to paint the same room how long would it take if smith and wesson worked together?

OpenStudy (anonymous):

okay thank you. also how would I write this as a rational equation?

OpenStudy (anonymous):

btw the answer is almost immediate \[\frac{4\times 5}{4+5}=\frac{20}{9}\]

OpenStudy (anonymous):

okay well you answered my question.

OpenStudy (anonymous):

the equation would be \[(\frac{1}{4}+\frac{1}{5})T=1\] where T is time

OpenStudy (anonymous):

when you add you get \[\frac{9}{20}T=1\] making \[T=\frac{20}{9}\]

OpenStudy (anonymous):

Thank you so much @satellite73